OpenStudy (anonymous):

Hey guys I have a silly question but please just tell me how to do it! The direction is found by using the tangent function: 3 miles/7miles = Tan(teta) = teta=23degrees my question is I am doing the calculation but from where they got 23 degrees?

zepdrix (zepdrix):

No that's a good question :) So the relationship of the vertical and horizontal components and the angle gave you:\[\Large \frac{3}{7} \quad=\quad \tan \theta\] To solve for theta, we can rewrite this in terms of the inverse tangent function by swapping the 3/7 and theta.\[\Large \theta=\arctan\left(\frac{3}{7}\right)\]

zepdrix (zepdrix):

I don't think you can find this angle by any simple refrence. You'd want to punch it into your calculator.

OpenStudy (phi):

if tan(A)= 3/7 then A = atan(3/7) calculator needed

zepdrix (zepdrix):

Oh, and make sure your calculator is in degree mode! :)

OpenStudy (anonymous):

you guys are awesome! thanks so much!